OpenStudy (factor):

2 numbers are in the ratio 15:11.If their hcf is 13 and lcm is 2145 then find the numbers

OpenStudy (factor):

@hartnn help

OpenStudy (factor):

ur gud in maths rite

hartnn (hartnn):

ok. The product of 2 numbers is also the product of its LCM and HCF so if the 2 numbers are x and y we have xy = 13*2145 got this ?

OpenStudy (factor):

cmon hartnn pls mntion the steps

OpenStudy (factor):

pls ellaborate @hartnn

hartnn (hartnn):

what did you not understand in that ? wasn't that very clear ?

OpenStudy (factor):

no how do

OpenStudy (factor):

i do it pls say it step by step

hartnn (hartnn):

that was 1st step : since xy = lcm * hcf xy = 13*2145

OpenStudy (factor):

okay

OpenStudy (factor):

secind step

hartnn (hartnn):

x/y = 15/11 so, x= 15y/11 plug this x in your first step equation!

OpenStudy (anonymous):

let the numbers be 15 x and 11 x ,then 15x*11x=13*2145 find x and then the numbers.